OpenStudy (anonymous):

is 0.9999999 the same as 1.0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if its repeated

OpenStudy (anonymous):

You can show it with a delta epsilon proof

OpenStudy (anonymous):

Given that there are infinite number of 9s the limiting value is 1

OpenStudy (anonymous):

my give medal button is gone <.<

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

Since there are an infinite number of 9s in the decimal expansion you must treat it as the limiting value as the number of 9s approaches infinity.

OpenStudy (anonymous):

yeah it is there are quite a few ways to see it, you can do an infinte geometric series: \[.9999999...=\frac{9}{10} + \frac{9}{100} +\frac{9}{1000} +\cdots = \frac{\frac{9}{10}}{1-\frac{1}{10}} = 1\]

OpenStudy (anonymous):

or you can try to find a number in between them, but you wont be able to: \[\frac{1+ .9999\ldots}{2} = .99999\ldots \]

OpenStudy (anonymous):

This all boils down to the concept of limit. Read about the delta epsilon definition of the limit.

OpenStudy (anonymous):

Let n = "number of nines" we want to show that for any e there is an N such that for all n > N 0.(n nines) is less than |1-e|

OpenStudy (anonymous):

That's easy to show, but my formulation is a bit messy.

OpenStudy (anonymous):

err |1 - 0.(n nines)|

OpenStudy (anonymous):

You get the idea.